# Spontaneous Symmetry Breaking of an Optical Polarization State in a Polarization-Selective Nonlinear Resonator

### v1

@@ Acknowledgments.
.1 Self rotation: experiment
@@
A scheme of the experimental setup for measuring PSR is shown in Fig. 5. The beam from an external cavity-diode laser (Vitawave ECDL-7950R) delivers approximately $50\mathrm{~{}mW}$ of optical power in a single longitudinal mode. The laser is tuned across the rubidium D1 line ($\lambda\simeq 795\mathrm{~{}nm}$). The beam is shaped by a pair of lenses and then transmitted through a Glan prism that transmits strictly linear polarization. A wave plate is placed before the Glan prism to vary the pump intensity; the transmitted intensity is 13.5 mW unless stated otherwise. The horizontally polarised light beam is then focused by a lens into a cylindrical vapour cell ($25\mathrm{~{}mm}$ in diameter and $75\mathrm{~{}mm}$ in length) with antireflection coated windows, filled with ${}^{85}\mathrm{Rb}$ without buffer gas at 70C. With an estimated focal beam diameter of $d=50\,\mu\mathrm{m}$, we experimentally obtain $I_{\rm sat}\simeq 0.8\mathrm{~{}mW}/\pi d^{2}$, thus operating in a high saturation regime.
To produce and control ellipticity in the initially horizontally polarised beam, a fine-adjustable $\lambda/4$ plate is placed in front of the cell. To compensate for the impact of birefringent windows on pumping light, we align the major axis of the polarisation ellipse with the optical axis of the cell windows. To this end, we place a $\lambda/2$ wave plate after the $\lambda/4$ plate and adjust it every time we rotate the latter to and check that the polarization pattern does not rotate when the laser is tuned into and out of the atomic line.
After exiting the cell, the beam is directed to the measuring system consisting of a $\lambda/2$ wave plate and a PBS, whose axis is aligned with the Glan prism. By rotating this $\lambda/2$ plate, we can measure the signal in either canonical or diagonal basis. The outputs of the PBS are directed to two photodiodes. Data are gathered by tuning the laser through the rubidium D1 line for a range of initial ellipticities and recording the signal from the photodiodes. The rotation angle of the polarisation ellipse is determined by (see Supplementary C)
@@
.2 Self-rotation in an atomic X system
@@
We develop the theory of PSR and Eq. (6) assuming X-shaped level structure as shown in Fig. 7. The Hamiltonian of this system under the rotating-wave approximation is as follows:
@@
$\dot{\rho}_{24}=$ $-\Omega_{R}^{*}\rho_{14}-\Omega_{L}\rho_{23}\overset{!}{=}0;$
$\dot{\rho}_{23}=$ $i\Delta\rho_{23}+\Omega_{R}^{*}\rho_{13}-\Omega_{L}^{*}\rho_{24}\overset{!}{=}0;$
$\dot{\rho}_{14}=$ $-i\Delta\rho_{14}+\Omega_{R}\rho_{24}-\Omega_{L}\rho_{13}\overset{!}{=}0;$
$\dot{\rho}_{13}=$ $-\Omega_{R}\rho_{23}-\Omega_{L}^{*}\rho_{14}\overset{!}{=}0$
@@
where we used Eq. (16) and that $|H|^{2}+|V|^{2}=|\mathcal{E}_{+}|^{2}+|\mathcal{E}_{-}|^{2}$.
To derive Eq. (7), we use the circular basis:
@@
.4 Full experimental setup
We use an additional rubidium cell as a reference to determine the laser frequency using absorption spectroscopy. Waveplates before and after the rubidium cell are needed to compensate the birefringence of the cell’s windows. Resonator losses are controlled by rotating the $\lambda/2$ plate in front of PBS3 in the resonator loop (Fig. 8).
To compensate for the birefringence of the cell windows and other imperfections of the setup, the polarization of the input beam is fine-tuned by two waveplates before the cell to eliminate the cavity resonance peaks when the laser is detuned from the atomic line. The cavity is aligned to maximize the mode matching between the horizontally polarized light returning from PBS3 and the horizontal pump by observing interference between the two fields. We achieve the maximum visibility of $V=0.96\pm 0.01$, which corresponds to a mode-matching efficiency of $V^{2}\approx 0.92$.

### v2

@@ Acknowledgments
I Self rotation: experiment
@@
A scheme of the experimental setup for measuring PSR is shown in Fig. 5. The beam from an external cavity-diode laser (Vitawave ECDL-7950R) delivers approximately $50\mathrm{~{}mW}$ of optical power in a single longitudinal mode. The laser is tuned across the rubidium D1 line ($\lambda\simeq 795\mathrm{~{}nm}$). The beam is shaped by a pair of lenses and then transmitted through a Glan prism that transmits strictly linear polarization. A wave plate is placed before the Glan prism to vary the pump intensity; the transmitted intensity is 13.5 mW unless stated otherwise. The horizontally polarised light beam is then focused by a lens into a cylindrical vapor cell ($25\mathrm{~{}mm}$ in diameter and $75\mathrm{~{}mm}$ in length) with antireflection coated windows, filled with ${}^{85}\mathrm{Rb}$ without buffer gas at 70C. With an estimated focal beam diameter of $d=50\,\mu\mathrm{m}$, we experimentally obtain $I_{\rm sat}\simeq 0.8\mathrm{~{}mW}/\pi d^{2}$, thus operating in a high saturation regime.
To produce and control ellipticity in the initially horizontally polarised beam, a fine-adjustable $\lambda/4$ plate is placed in front of the cell. To compensate for the impact of birefringent windows on pumping light we used a procedure described in Section 4.
After exiting the cell, the beam is directed to the measuring system consisting of a $\lambda/2$ wave plate and a PBS, whose axis is aligned with the Glan prism. By rotating this $\lambda/2$ plate, we can measure the signal in either a canonical or diagonal basis. The outputs of the PBS are directed to two photodiodes. Data are gathered by tuning the laser through the rubidium D1 line for a range of initial ellipticities and recording the signal from the photodiodes. The rotation angle of the polarisation ellipse is determined by (see Supplementary 3)
@@
II Self-rotation in an atomic X system
@@
We develop the theory of PSR and Eq.6 from the main text assuming X-shaped level structure as shown in Fig. 7. The Hamiltonian of this system under the rotating-wave approximation is as follows:
@@
$\dot{\rho}_{24}=$ $i\Omega_{R}^{*}\rho_{14}-i\Omega_{L}\rho_{23}\overset{!}{=}0;$
$\dot{\rho}_{23}=$ $i\Delta\rho_{23}+i\Omega_{R}^{*}\rho_{13}-i\Omega_{L}^{*}\rho_{24}\overset{!}{=}0;$
$\dot{\rho}_{14}=$ $-i\Delta\rho_{14}+i\Omega_{R}\rho_{24}-i\Omega_{L}\rho_{13}\overset{!}{=}0;$
$\dot{\rho}_{13}=$ $i\Omega_{R}\rho_{23}-i\Omega_{L}^{*}\rho_{14}\overset{!}{=}0$
@@
where we used Eq. 16 and that $|H|^{2}+|V|^{2}=|\mathcal{E}_{+}|^{2}+|\mathcal{E}_{-}|^{2}$.
To derive Eq. 7 from the main text, we use the circular basis:
@@
IV Full experimental setup
@@
We use an additional rubidium cell as a reference to determine the laser frequency using absorption spectroscopy.
Waveplates before and after the rubidium cell are needed to compensate for the birefringence of the cell’s windows. To compensate for the birefringence of the cell windows and other imperfections of the setup, the polarization of the input beam is fine-tuned by these waveplates to eliminate the cavity resonance peaks when the laser is detuned from the atomic line.
The cavity is aligned to maximize the mode matching between the horizontally polarized light returning from PBS3 and the horizontal pump by observing interference between the two fields. We achieve the maximum visibility of $V=0.96\pm 0.01$, which corresponds to a mode-matching efficiency of $V^{2}\approx 0.92$.
Resonator losses are controlled by rotating the $\lambda/2$ plate between PBS1 and PBS3 in the resonator loop (Fig. 8).